# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

### v2

@@ Abstract
Identifying variants that carry substantial information on the trait of interest remains a core topic in genetic studies. In analyzing the EADB-UKBB dataset to identify genetic variants associated with Alzheimer’s disease (AD), however, we recognize that both existing marginal association tests and conditional independence tests using knockoffs suffer either power loss or lack of informativeness, especially when strong correlations exist among variants. To address the limitations of existing knockoff filters, we propose a new feature-versus-group (FVG) filter that is more powerful and precise in identifying important features from a set of strongly correlated features using group knockoffs. In extensive simulation studies, the FVG filter controls the expected proportion of false discoveries and identifies important features with enhanced power and greater precision. Applying the proposed method to the EADB-UKBB dataset, we discover important variants from 84 loci (same as the most powerful group knockoff filter) with catching sets of substantially smaller size and higher purity.
Keywords: Alzheimer’s disease genetics, False discovery rate, Feature selection, Knockoffs.
@@ 1.1 Conditional Independence Test to Discover Disease-associated Variants
Testing conditional independence between a set of features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the response of interest $Y$ is an important topic in various research areas, including causal inference (Peters,, 2015; Cai et al.,, 2022), genetic analysis (Khera and Kathiresan,, 2017; Zhu et al.,, 2018) and graphical model learning (Deka et al.,, 2016; Tugnait,, 2022). In the era of big data, as the number of features increases, the need for statistical approaches for simultaneous inference of conditional independence between hundreds of thousands of features and the response keeps emerging.
@@
Specifically, developing statistical methods for conditional independence tests with type-I error rate control has been an important and popular research topic of large-scale genome-wide association studies (GWAS; Tang and He,, 2021; Hou et al.,, 2023; Morra et al.,, 2023) in recent years. With the ultimate goal of identifying novel targets for the development of genomic-driven medicine, the contribution of more genetic variants to complex phenotypes has been investigated in GWAS. However, our ability to detect causal genetic variants and to translate discoveries into insights of the underlying genetic mechanisms does not increase proportionally to the growth of the scale of GWAS. On one hand, with the stringent criterion to control the family-wise error rate (FWER), conventional GWAS are usually suboptimal in terms of the statistical power of detecting important variants. On the other hand, because conventional GWAS base their inference on marginal association models which regress $Y$ on one $X_{j}$ at a time, they can only provide blurred results with a lot of proxy variants correlated with the true causal variants (Schaid et al.,, 2018). Both issues become more severe when more variants are sequenced at a higher resolution because the $p$-value threshold of FWER control also becomes more stringent and there are more proxy variants strongly correlated with true causal variants. Although many fine-mapping methods have been proposed (Benner et al.,, 2017; Zou et al.,, 2022) to pinpoint the causal variants, they can only partially address the aforementioned issues without control of type-I error rate, calling for new statistical inference methods with better theoretical properties and empirical performance in multiple testing of conditional independence for real-world genetic data analysis.
@@
In this article, we investigate the stage I meta-analysis of the European Alzheimer & Dementia Biobank (EADB) dataset and the UK Biobank (UKBB) dataset (Bellenguez et al.,, 2022) to identify genetic variants associated with AD. This meta-analyzed EADB–UKBB dataset is deposited in the European Bioinformatics Institute GWAS Catalog [^1] https://www.ebi.ac.uk/gwas/ (Accession number: GCST90027158), including $39,106$ clinically diagnosed AD cases, $46,828$ proxy cases, and $401,577$ control cases from $15$ European countries. We focus on $640,576$ directly genotyped variants with minor allele frequency (MAF) at least $0.01$, resulting in a processed dataset $\mathbb{G}$ of $487,511$ observations and $640,576$ features. With such a large dataset, we aim to discover as many important genetic variants that carry independent information of AD without making too many false discoveries. Mathematically, such a task can be transferred into a multiple testing problem of conditional independence between a large set of variants $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the binary response $Y$ which indicates whether a case is clinically diagnosed AD.
@@ 1.2 Existing Approaches and Their Limitations
Recent decades have witnessed fruitful works in the development of multiple testing procedures. To control the family-wise error rate (FWER) of making at least one false discovery, a lot of $p$-value-based methods have been developed as improvements of the classic Bonferroni correction, including the Šidák correction (Šidák,, 1967), Holm’s step-down procedure (Holm,, 1979) and Hochberg’s step-up procedure (Hochberg,, 1988). However, FWER has been criticized for its conservativeness in many settings, especially when most of the signals are weak. To perform powerful inference, Benjamini and Hochberg, (1995) proposed the false discovery rate (the expected proportion of false discoveries in all discoveries) as an alternative type-I error rate measure and developed a procedure to perform multiple testing with provable FDR control under the assumption that $p$-values are independent or positively dependent. Inspired by this pioneering work, Benjamini and Yekutieli, (2001) developed a generalized procedure without assumptions on dependencies among $p$-values, while Storey, (2002) and Whittemore, (2007) investigated the Bayesian counterpart of FDR. However, as most of the aforementioned methods require $p$-values of nulls for all features, their feasibility on multiple testing of conditional independence becomes questionable when the number of features $p$ greatly exceeds the sample size $n$. The main reason is that $p$-values of conditional independence under high-dimensional scenarios are nontrivial and generally not available except in a very few specific cases (Tibshirani et al.,, 2016; Lee et al.,, 2016).
Unlike all the above methods, the recently developed knockoff filter (Barber and Candès,, 2015; Candès et al.,, 2018) and its variants (Gimenez et al.,, 2019; Gimenez and Zou,, 2019; Barber and Candès,, 2019; Huang and Janson,, 2020; Ren and Candès,, 2023) are a family of approaches that can control finite-sample FDR in multiple testing of conditional independence hypotheses at the feature-versus-feature level
@@
without any assumption on the conditional distribution $Y|\textbf{X}$. Specifically, these methods first construct knockoffs $\widetilde{\textbf{X}}$ that mimic the dependency among features X and are not dependent to the response $Y$. By doing so, knockoffs play the role of synthetic controls: if $H^{(\text{ff})}_{j}$ is true, ${X}_{j}$ and $\widetilde{X}_{j}$ are indistinguishable even with information of $Y$; if $H^{(\text{ff})}_{j}$ is false, ${X}_{j}$ and $\widetilde{X}_{j}$ possess different dependencies with $Y$. By comparing dependencies between $Y$ and X with dependencies between $Y$ and $\widetilde{\textbf{X}}$, these methods have both controlled FDR and superior power in detecting true causal features as validated in extensive simulation studies.
However, in the analysis of real-world genetic data, methods that infer $H^{(\text{ff})}_{j}$’s usually suffer severe power loss. The main reason is that linkage disequilibrium could introduce high correlations among genetic variants, making it difficult to distinguish variants that are truly associated with the response (whose $H^{(\text{ff})}_{j}$’s are false) from their strongly correlated null variants (whose $H^{(\text{ff})}_{j}$’s are true). This would significantly decrease the number of discoveries as shown by the simulation studies in Barber and Candès, (2015) and Candès et al., (2018). Such a power loss issue would deteriorate when more genetic variants with strong correlations are recorded under a growing resolution of sequencing. For example, when we apply the model-X knockoff filter (Candès et al.,, 2018) to the EADB-UKBB dataset (Bellenguez et al.,, 2022), only 138 AD-associated genetic variants from 49 loci (here, we define two loci as different if they are at least 1Mb away from each other) are identified under the target FDR level $0.10$ as shown in Table 1. Compared with the 51 loci discovered by marginal association test ($p$-value threshold: $5\times 10^{-8}$) in Bellenguez et al., (2022), although model-X knockoff filter identified 6 new loci, it misses several important AD loci with strong signals, including the “FAM193B” locus, the “FAM157C” locus, the “NSF” locus, the “SIGLEC11” locus and the “APP” locus.
Table 1: Summary of results by applying different methods to the EADB-UKBB dataset.
@@
Marginal association test 51 18.078 18.078 0.477
Model-X knockoff filter 49 2.816 1.000 1.000
Group knockoff filter 84 23.714 13.105 0.610
@@
To address this power loss issue, Dai and Barber, (2016) proposed the group knockoff filter. By defining groups $B_{1},\ldots,B_{k}$ such that features with strong correlations are in the same group and changing the inference target as
@@
at the group-versus-group level, important features and their strongly correlated null features can be rejected as a group. Because the grouping is done to ensure features from different groups are weakly correlated, the group knockoff filter can circumvent the obstacle of high correlation in inferring $H^{(\text{ff})}_{j}$ with significantly higher power. To apply the group knockoff filter to the EADB-UKBB dataset (Bellenguez et al.,, 2022), we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct $321,535$ variant groups by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, variants in the EADB-UKBB dataset (Bellenguez et al.,, 2022) are grouped such that the average absolute correlation of variants from different groups is lower than 0.5. Under such a group structure, the group knockoff filter identifies 152 AD-associated variant groups from 84 loci under the target FDR level $0.10$. Specifically, among these 152 identified groups, 126 are located outside the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650), including groups in loci “FAM193B”, “FAM157C”, “NSF”, “SIGLEC11” and “APP” that are missed by the model-X knockoff filter. In addition, the group knockoff filter also identifies 39 new loci that are not discovered by Bellenguez et al., (2022).
@@
However, doing so can only provide a series of AD-associated variant groups without information on which variants contribute more to AD variation within each group, leaving the target of identifying AD-associated variants not fully addressed. This can be observed from Table 1. With an average size $13.105$ of the identified groups given by the group knockoff filter, 35 identified groups contain at least 10 genetic variants and the 5 largest groups contain 314, 232, 136, 121 and 117 variants respectively. With such large sizes, which genetic variants in these groups contribute more to AD variation remains unanswered.

### v3

# Refining Genetic Discoveries of Group Knockoffs via A Feature-level Filter

@@ Abstract
Identifying variants that carry substantial information on the trait of interest remains a core topic in genetic studies. In analyzing the EADB-UKBB dataset to identify genetic variants associated with Alzheimer’s disease (AD), however, we recognize that both existing marginal association tests and conditional independence tests using existing knockoff filters suffer either power loss or lack of informativeness, especially when strong correlations exist among variants. To address these limitations, we propose a new feature-versus-group (FVG) filter that achieves balance between the power and precision in identifying important features from a set of strongly correlated features using group knockoffs. In extensive simulation studies, the FVG filter controls the expected proportion of false discoveries and identifies important features in smaller catching sets without large power loss. Applying the proposed method to the EADB-UKBB dataset, we discover important variants from 89 loci (similar to the most powerful group knockoff filter) with catching sets of substantially smaller size and higher purity and verify the biological informativeness of our discoveries.
Keywords: Alzheimer’s disease genetics, False discovery rate, Genetic variant selection, Knockoffs.
@@ 1.1 Conditional Independence Test to Discover Disease-associated Variants
Testing conditional independence between a set of features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the response of interest $Y$ is an important topic in various research areas, including causal inference (Peters2015; Cai2022), genetic analysis (Khera2017; Zhu2018) and graphical model learning (Deka2016; Tugnait2022). In the era of big data, as the number of features increases, the need for statistical approaches for simultaneous inference of conditional independence between hundreds of thousands of features and the response keeps emerging.
@@
Specifically, developing statistical methods for conditional independence tests with type-I error rate control has been an important and popular research topic of large-scale genome-wide association studies (GWAS; Tang2021; Hou2023; Morra2023) in recent years. With the ultimate goal of identifying novel targets for the development of genomic-driven medicine, the contribution of more genetic variants to complex phenotypes has been investigated in GWAS. However, our ability to detect causal genetic variants and to translate discoveries into insights of the underlying genetic mechanisms does not increase proportionally to the growth of the scale of GWAS. On one hand, with the stringent criterion to control the family-wise error rate (FWER), conventional GWAS are usually suboptimal in terms of the statistical power of detecting important variants. On the other hand, because conventional GWAS base their inference on marginal association models which regress $Y$ on one $X_{j}$ at a time, they can only provide blurred results with a lot of proxy variants correlated with the true causal variants (Schaid2018). Both issues become more severe when more variants are sequenced at a higher resolution because the $p$-value threshold of FWER control also becomes more stringent and there are more proxy variants strongly correlated with true causal variants. This calls for new statistical inference methods with better theoretical properties and empirical performance in multiple testing of conditional independence for real-world genetic data analysis.
@@
In this article, we investigate the stage I meta-analysis of the European Alzheimer & Dementia Biobank (EADB) dataset and the UK Biobank (UKBB) dataset (Bellenguez2022) to identify genetic variants associated with AD. This meta-analyzed EADB–UKBB dataset is deposited in the European Bioinformatics Institute GWAS Catalog (https://www.ebi.ac.uk/gwas/; accession number: GCST90027158), including $39,106$ clinically diagnosed AD cases, $46,828$ proxy cases, and $401,577$ control cases from $15$ European countries. We focus on $650,706$ directly genotyped variants with minor allele frequency (MAF) at least $0.01$, resulting in a processed dataset $\mathbb{G}$ of $487,511$ observations and $650,706$ features. With such a large dataset, we aim to discover as many as possible important genetic variants that carry independent information of AD without making too many false discoveries. Mathematically, such a task can be transferred into a multiple testing problem of conditional independence between a large set of variants $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the binary response $Y$ which indicates whether a case is clinically diagnosed AD.
@@ 1.2 Existing Approaches and Their Limitations
Recent decades have witnessed fruitful works in the development of multiple testing procedures. To control the family-wise error rate (FWER) of making at least one false discovery, a lot of $p$-value based methods have been developed as improvements of the classic Bonferroni correction, including the Šidák correction (Sidak1967), Holm’s step-down procedure (Holm1979) and Hochberg’s step-up procedure (Hochberg1988). However, FWER has been criticized for its conservativeness in many settings, especially when most of the signals are weak. To perform powerful inference, Benjamini1995 proposed the false discovery rate (the expected proportion of false discoveries in all discoveries) as an alternative type-I error rate measure and developed a procedure to perform multiple testing with provable FDR control under the assumption that $p$-values are independent or positively dependent (Yekutieli2001). Inspired by this pioneering work, Yekutieli2001 developed a generalized procedure without assumptions on dependencies among $p$-values, while Storey2002 and Whittemore2007 investigated the Bayesian counterpart of FDR. However, as most of the aforementioned methods require $p$-values of nulls for all features, their feasibility on multiple testing of conditional independence becomes questionable when the number of features $p$ greatly exceeds the sample size $n$. The main reason is that $p$-values of conditional independence under high-dimensional scenarios are nontrivial and generally not available except in a very few specific cases (Tibshirani2016; Lee2016).
Unlike all the above methods, the recently developed model-X knockoff filter (Candes2018) and its variations (Gimenez2019; Gimenez2019b; Barber2019; Bates2020; Huang2020; Ren2023; Li2024) are a family of approaches that can control finite-sample FDR in multiple testing of conditional independence hypotheses at the feature-versus-feature level
@@
without any assumption on the conditional distribution $Y|\textbf{X}$. Specifically, these methods first construct knockoffs $\widetilde{\textbf{X}}$ that mimic the dependency among features X and are not dependent on the response $Y$. By doing so, knockoffs play the role of synthetic controls: if $H^{(\text{ff})}_{j}$ is true, ${X}_{j}$ and $\widetilde{X}_{j}$ are indistinguishable even with information of $Y$; if $H^{(\text{ff})}_{j}$ is false, ${X}_{j}$ and $\widetilde{X}_{j}$ possess different dependencies with $Y$. By comparing the dependencies between $Y$ and X with those between $Y$ and $\widetilde{\textbf{X}}$, these methods have both controlled FDR and superior power in detecting true causal features as validated in extensive simulation studies.
However, in the analysis of real-world genetic data where genetic variants are the features to be selected, methods that infer $H^{(\text{ff})}_{j}$’s usually suffer severe power loss. The main reason is that linkage disequilibrium (LD) could introduce high correlations among genetic variants, making it difficult to distinguish variants that are truly associated with the response (whose $H^{(\text{ff})}_{j}$’s are false) from their strongly correlated null variants (whose $H^{(\text{ff})}_{j}$’s are true). This would significantly decrease the number of discoveries as shown by the simulation studies in Barber2015 and Candes2018. Such a power loss issue would deteriorate when more genetic variants with strong correlations are recorded under a growing resolution of sequencing. For example, when we apply the model-X knockoff filter (Candes2018) to the EADB-UKBB dataset (Bellenguez2022) , only 162 AD-associated genetic variants from 38 loci (here, we define two loci as different if they are at least 1Mb away from each other) are identified under the target FDR level $0.10$ as shown in Table 1. Compared with the 54 loci discovered by marginal association test ($p$-value threshold: $5\times 10^{-8}$) in Bellenguez2022, although model-X knockoff filter identified 7 new loci, it misses several important AD loci with strong signals, including the “ZCWPW1” locus on chromosome 7, the “PTK2B” locus on chromosome 8, the “ECHDC3” locus on chromosome 10, the ”MS4A4E” locus on chromosome 11, the “FERMT2” locus on chromosome 14, the “FAM157C” locus on chromosome 16 and the “SIGLEC11” locus on chromosome 19. This can seen in Figures 1 (a)-(b) where we displays the difference of inference results between marginal association test and model-X knockoff filter within a high LD region between positions 31571218 $\sim$ 32682663, chromosome 6. As correlation among genetic variants is high in this region, most of genetic variants possess a small $p$-values, which not only implies the existence of contributing variants but also obstacles our distinguishing of the contributing ones from all proxy ones. Such great correlations also result in no discoveries (under FDR $0.10$) of contributing variants of the model-X knockoff filter as the importance score of the most contributing variant does not exceed the data-driven threshold.
Table 1: Summary of results by applying different existing methods to the EADB-UKBB dataset, where bold values indicate the best performance under different evaluation metrics.
@@
Marginal association test 54 17.093 17.093 0.489
Model-X knockoff filter 38 4.263 1.000 1.000
Group knockoff filter 91 19.923 9.157 0.651
KnockoffZoom (2 layers) 91 15.152 5.621 0.851
KnockoffZoom (3 layers) 91 14.681 5.409 0.858
KeLP (2 layers) 45 10.422 2.535 0.918
KeLP (3 layers) 38 10.605 2.488 0.932
@@
(d) resolution-adaptive discoveries made by KnockoffZoom and KeLP that can be either variants identified
   at the variant level or variant groups identified at the group level containing no identified variants.
@@
: Size and purity of catching sets provided by model-X knockoff filter are trivially 1 and thus not included
   in comparison.
Refer to captionRefer to caption
(a) Manhattan plot of marginal association test and the correlation plot within the high-LD region.
(b) Manhattan plot of feature importance statistics of
different genetic variants under model-X knockoff filter.
(c) Manhattan plot of the contribution of different variants to the importance statistics of the group under the group knockoff filter.
Figure 1: Comparison of inference results provided by marginal association test (Bellenguez2022), model-X knockoff filter (Candes2018) and the group knockoff filter (Dai2016) in a high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6. (Displaying order of genetic variants is reorganized according to the hierarchical clustering of $(1-|\text{cor}(G_{i},G_{j})|)$ for better visualization.)
To address this power loss issue in real-world genetic data analysis, people turn their inference target from the feature-versus-feature level to the group-versus-group level and use group knockoff filter to detect signals lie in high LD region. First introduced by Dai2016 and developed by multiple works (Katsevich2019; Sesia2020; Sesia2021; Spector2022; Chu2024; Gablenz2024), the group knockoff filter defines groups $B_{1},\ldots,B_{k}$ such that features with strong correlations are empirically allocated in the same group and perform multiple testing on
@@
at the group-versus-group level, where important features and their strongly correlated null features can be rejected as a group. Because the grouping is done to ensure features from different groups are weakly correlated, the group knockoff filter can circumvent the obstacle of high correlation in inferring $H^{(\text{ff})}_{j}$ with significantly higher power. To apply the group knockoff filter to the EADB-UKBB dataset (Bellenguez2022), we generate the second-order group knockoffs as follows by assuming that the distribution of X can be well approximated by a multivariate Gaussian distribution. To empirically construct the genetic variant groups according to the correlation, we first use the UK Biobank directly genotyped data (which also records genetic information of Europeans) as the reference panel and compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ via the Pan-UKB consortium (https://pan.ukbb.broadinstitute.org). We then perform the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$ and construct $321,569$ variant groups. By doing so, the average absolute correlation of variants (features) from different groups is lower than 0.5. With second-order group knockoffs generated under such a group structure and the maximum-entropy construction using Algorithm 2 of Chu2024, the group knockoff filter identifies 198 AD-associated variant groups from 91 loci under the target FDR level $0.10$. Specifically, among these 198 identified groups, 121 are located outside the strongest APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650), including groups in loci “ZCWPW1”, “PTK2B”, “ECHDC3”, ”MS4A4E”, “FERMT2”, “FAM157C” and“SIGLEC11” that are missed by the model-X knockoff filter. In addition, the group knockoff filter also identifies 43 new loci that are not discovered by Bellenguez2022.
@@
However, doing so can only provide a series of AD-associated variant groups without information on which variants contribute more to AD variation within each group, leaving the target of identifying AD-associated variants not fully addressed. This can be observed from Table 1, where each identified groups given by the group knockoff filter are treated as a catching set of genetic signals. With an average size $19.923$ of catching sets, 30 catching sets contain at least 10 genetic variants and the 5 largest sets contain 314, 232, 117, 84 and 57 variants respectively. With such large sizes, which genetic variants in these catching sets contribute more to AD variation remains unanswered. This can also be seen in Figure 1 (c). Within the high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6, all genetic variants are clustered in the same group and such a group is identified as a sole catching set. However, such an inference result loses the information that all of the importance of this variant group is contributed by the variant rs9469112 as shown in Figure 1 (c).
To improve the precision of catching sets of genetic signals, several approaches have been developed to return catching sets of multiple resolutions or in a resolution-adaptive manner. One of the pioneering method is the multilayer knockoff filter (Katsevich2019) which simultaneously infers $H^{(\text{ff})}_{j}$’s and $H^{(\text{gg})}_{k}$’s, while the later developed KnockoffZoom [^1] In this article, we adopt Algorithm 5 in the supplementary note 4 of Sesia2020. (Sesia2020) and knockoff e-value linear program (KeLP; Gablenz2024) are implemented to identify genetic signals in resolution-adaptive manners by using catching sets at different resolutions. Specifically, KnockoffZoom gradually refines catching sets by sequentially performing group knockoff inference from lower resolutions to higher resolutions and finally using model-X knockoff filter at the feature-versus-feature level. KeLP, in contrast, utilizes the method of Ren2024 to compute $e$-values of all variant groups at different resolutions and all variants and identifies signals at the finest possible resolution by solving an optimization problem with nesting constraints.
However, when we apply KnockoffZoom and KeLP to identify contributing genetic variants under the target FDR level $0.10$ at both the feature-versus-feature level as in the model-X knockoff filter and group-versus-group level as in the group knockoff filter (which we label as “(2 layers)”), existing issues are not fully addressed. On one hand, as the inference of KnockoffZoom (2 layers) at the group-versus-group level is the same as the group knockoff filter, KnockoffZoom (2 layers) identifies the same loci as the group knockoff filter and does not suffer from any power loss. However, as signal-to-noise ratios of $H^{(\text{ff})}_{j}$’s are low due to high LD, half of groups (99/198) contain no identified variants at the feature-versus-feature level. Thus, KnockoffZoom (2 layers) only records moderate improvement of average size (decreases from 9.157 to 5.621) and purity (increases from 0.651 to 0.851) of catching sets. This can be seen from Figure 7 (a) in Appendix A, where importance score of the most contributing variant in the high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6, remains lower than the data-driven threshold. Even when we add an intermediate layer (obtained by the hierarchical clustering with average linkage and cutoff value $0.25$) between the current two levels in applying KnockoffZoom (we refer to as “KnockoffZoom (3 layers)”), such a phenomenon retains as shown in Figures 7 (b)-(c). KeLP, on the other hand, suffer from greater power loss even compared to the model-X knockoff filter, no matter with 2 layers or 3 layers. The main reason is that computing e-values for resolution-adaptive inference under the target FDR level $0.10$ requires inferences under a target FDR level lower than $0.10$ throughout all resolutions. As only variants and groups with nonzero e-values can be selected by KeLP, lower target FDR levels would decrease the number of variants and groups with nonzero e-values and thus the number of catching sets.
Another category of commonly used approaches are various Bayesian fine-mapping approaches (Wilson2010; Guan2011; Hormozdiari2014; Chen2015; Benner2016; Wang2020; Zou2022; Li2024a). For comprehensive review, please refer to Schaid2018 and Li2025. Generally, existing Bayesian fine-mapping approaches fit a parametric model between the response and features (genetic variants) in identified important regions, compute the posterior inclusion probability (PIP) of each feature to be contributive via Bayesian approaches and construct credible sets of features to contain at least one signal with large posterior probability. Potentially, one can perform fine-mapping within variant groups identified by the group knockoff filter and obtain credible sets of smaller size and higher precision. However, using such a two-step heuristic approach still faces the tradeoff between power and precision via the selection of coverage probability threshold.